# MambaFoley: Foley Sound Generation using Selective State-Space Models

### v1

@@ Abstract
Recent advancements in deep learning have led to widespread use of techniques for audio content generation, notably employing Denoising Diffusion Probabilistic Models (DDPM) across various tasks. Among these, Foley Sound Synthesis is of particular interest for its role in applications for the creation of multimedia content. Given the temporal-dependent nature of sound, it is crucial to design generative models that can effectively handle the sequential modeling of audio samples. Selective State Space Models (SSMs) have recently been proposed as a valid alternative to previously proposed techniques, demonstrating competitive performance with lower computational complexity. In this paper, we introduce MambaFoley, a diffusion-based model that, to the best of our knowledge, is the first to leverage the recently proposed SSM known as Mamba for the Foley sound generation task. To evaluate the effectiveness of the proposed method, we compare it with a state-of-the-art Foley sound generative model using both objective and subjective analyses.
@@
Foley Sound Synthesis, State Space Models, Generative Sound Synthesis, Denoising Diffusion Probabilistic Models
@@ I Introduction
In recent years, audio generation using deep-learning generative techniques has made relevant progress. This was possible primarily by the adoption of generative models denoted as Denoising Diffusion Probabilistic Models (DDPMs) [1], which have been applied to a wide variety of audio application such as Text-to-Audio generation [2, 3, 4], sound field reconstruction [5], among others. Generated audio samples from generative models are so realistic that using them to replace real datasets for training deep learning models has been explored across various domains [6, 7].
@@
Generating audio waveforms with mid and long-term temporal coherency has always been a challenge for learning-based models. In fact, although using raw waveforms in audio generative modeling can significantly improve the quality of generated audio content, they also pose challenges due to the high-dimensional input signal, which includes tens of thousands of samples. Several deep learning models have been proposed to address the challenge of modeling raw waveform sequences, each with its own drawbacks. RNNs based on LSTMs [17] are not parallelizable, which limits their computational efficiency. CNNs [18] are constrained by their receptive field, which limits their ability to model the global coherence of the waveform. Transformers and in general attention-based methods [19, 20], while effective for modeling sequential data, are impractical for very long sequences due to their quadratic complexity with respect to window length. Recently, structured State-Space Models (SSMs) [21, 22] have been proposed as a new class of architectures tailored to sequence modeling, combining characteristics of both CNNs and RNNs. Architectures such as SaShiMi [23] and especially Mamba [24, 25] outperformed previous state-of-the-art techniques over waveform modeling while scaling linearly in complexity with respect to sequence length. As expected, Mamba models were readily applied to a diverse variety of audio signal processing tasks such as audio representation learning [26], deepfake detection [27], and speech separation [28]. Motivated by previous studies, we explored how to integrate selective SSMs into the Foley sound generation task.
In this paper, we introduce MambaFoley, a method for Foley sound generation that leverages the SSM Mamba to model raw audio. To the best of our knowledge, this is the first application of Mamba in Foley sound generation. The proposed model employs the diffusion architecture DAG [29] and integrates Mamba in the bottleneck to improve its sequence modeling capabilities. Additionally, we enable temporal control of the generated Foley sounds through the BlockFiLM conditioning method introduced in [14]. Fig. 1 shows a schematic overview of the MambaFoley model. To validate the effectiveness of our proposed model, we conduct a series of experiments comparing it with the model proposed in [14] and an Attention-based baseline. We evaluate performance using both objective metrics and subjective evaluations. The rest of the paper is organized as follows: Sec. II provides the reader with the necessary background on SSMs. In Sec. III, we present design choices of the MambaFoley model. Sec. IV describes how the experiments were organized, while Sec. V presents the results, highlighting the effectiveness of the proposed technique in comparison to state-of-the-art methods. In Sec. VI we draw some conclusions. Code and listening examples are publicly available online [^1] https://github.com/FurioColombo/mamba-foley.
@@ III-A Problem formulation
The goal of MambaFoley is to generate raw audio waveforms that represent specific sound classes (e.g. dog, gunshot, …) while also following a specified temporal pattern (e.g. two dog barks separated by one second of silence).
More formally, let $\mathbf{x}\in\mathbb{R}^{N}$ be a discrete audio waveform whose content corresponds to the category indicated by class $c\in\{0,\ldots,C-1\}$ of length $N$ samples We want to control the temporal evolution of $\mathbf{x}$ using $\boldsymbol{\tau}\in\mathbb{R}^{I}$, which can be defined using different representations.
MambaFoley then learns the function $\mathcal{M}(\cdot)$ such that:
$\mathbf{x}=\mathcal{M}(c,\boldsymbol{\tau}).$ (6)
Following the approach proposed in [14], we condition the temporal evolution of the generated audio using the Root-Mean-Square (RMS) of a waveform signal that contains an audio excerpt of the target category with the desired temporal evolution. Specifically, in our scenario we can define:
@@ III-B Network Architecture
MambaFoley is a diffusion-based model, as such, the target audio waveform $\mathbf{x}$ is progressively corrupted with Gaussian noise throughout the forward process. In the backward process, the original waveform is iteratively denoised using a U-Net architecture, whose parameters are updated based on the $\ell_{2}$ norm between the predicted and ground truth noise $\boldsymbol{\epsilon}$. The overall architecture of MambaFoley is illustrated in Fig. 2.
We select the DAG architecture [29] as the U-Net backbone of our model, due to its proven effectiveness in generating high-resolution raw audio. We introduced some changes to DAG, specifically regarding the conditioning and bottleneck of the model.
Similar to DAG, we refer to the building blocks of MambaFoley as GBlocks, to which we applied significant modifications. GBlocks consist of sequences of 1D-convolutional layers in the downsampling encoder part, or transposed 1D-convolutions in the upsampling decoder. Fig. 3 (a) shows the architecture of GBlocks.
@@
In the bottleneck of the model, between the encoder and decoder parts of the U-Net, we introduce a bidirectional Mamba layer [28] to exploit the capabilities of SSMs in processing temporal audio sequences. This layer consists of two parallel Mamba blocks, as defined in Sec. II, one processing the input sequence and the other one processing its reversed version. Fig. 3 (b) shows the bidirectional Mamba bottleneck. The outputs of the two parallel Mamba blocks are each normalized using Root Mean Square Layer Normalization (RMSNorm) [31], chosen for its computational efficiency. The normalized outputs are then summed with the input (and reversed) signal residuals. Finally, the outputs of the two SSMs are concatenated and processed by a single linear layer. We use state dimension D=$16$, expansion factor $4$ and local convolution width $4$ for the Mamba layers.
@@ IV-C Training details
All models are trained for $500$ epochs using the Adam [35] optimizer with a learning rate of 1e-4. To make the training more robust, we randomly train the network unconditionally with a $0.1$ probability. We adopt a variance-preserving cosine scheduler, as proposed in  [36]. For conditional sampling, we employ the DDPM-like discretization of SDE [37] with classifier-free guidance [38] to enhance sample diversity and quality. For the temporal event features, we use a window length of $W=512$ samples and a hop size of $H=128$ samples, which correspond to approximately $24~{}\mathrm{ms}$ and $6~{}\mathrm{ms}$, respectively, given the sampling rate of $22.05$ kHz. For a fair comparison with to T-Foley [14], we also exclude the SneezeCough class from the listening test.
@@
Figure 4: Example spectrograms obtained using the different methods. The category conditioning corresponds to the class gunshot, while the temporal conditioning is provided via the RMS shown in (a) and computed over the ground truth shown in (b).
@@ V Results
In this section, we present results demonstrating the effectiveness of the proposed MambaFoley model compared to T-Foley and AttentionFoley baselines, both in terms of objective and subjective evaluation metrics. Before presenting the analyses computed on the entire dataset, we first examine a representative spectrogram generated from the waveforms obtained through each of the considered models, shown in Fig. 4. The STFT for all spectrograms is computed with a Hann window of length $512$, hop size of $128$ samples and a FFT of $512$ points. Fig. 4(b) contains an example of the target sound, corresponding to a gun shot, from which we extract the RMS, shown in Fig.4 (a). The RMS is then used to condition the temporal evolution of the network. As observed from the spectrograms of all the considered models in Fig. 4(c) (MambaFoley), (d) (T-Foley), and (e) (AttentionFoley), each model appears to accurately reproduce the desired temporal conditioning. However, MambaFoley seems to reconstruct more content at higher frequencies.
@@ V-A Objective Evaluation Results
Real data - 0 19.11 15.20 3.31
T-Foley 70.299,745 0.0351 59.48 40.41 8.60
AttentionFoley 63.327,329 0.0393 60.94 40.58 8.35
MambaFoley 70.196,449 0.0374 58.65 35.43 7.77
Table I reports the results for E-L1 and FAD for the three different feature extractors considered in this study, along with the number of parameters for each network. To ease the analysis, in the top row we show the lower bound on the achievable error, namely the FAD and E-L1 scores computed on the real data.Although the lower bound of the E-L1 distance is clearly 0, the FAD calculated between the embeddings of the test and train datasets for real data is greater than zero. This indicates the inherent differences between real audio samples. This finding helps in estimating the lower bound for FAD in our task, where our goal is to generate audio samples that are as realistic as possible.
As it is possible to observe from Table I, MambaFoley outperforms both T-Foley and AttentionFoley in terms of FAD, while achieving comparable results in E-L1. This is probably due to the role of the conditioning blocks over the bottleneck, which ensures consistent temporal alignment. Nevertheless, these results further support the effectiveness of the Mamba block’s sequence modeling capabilities. It not only enhances the overall quality of the generated audio but also preserves temporal audio features.
@@ VI Conclusion
In this paper, we introduced MambaFoley, a Denoising Diffusion Probabilistic model for Foley sound generation exploiting the power of selective State Space models. Specifically, we have applied Mamba in the bottleneck of a diffusion model, a SSM that has proved to obtain notable performance in audio sequence modeling, while mantaining computational complexity. Through a series of experiments that include both objective and subjective results, we demonstrate the effectiveness of the proposed approach by comparing it with baseline methods using various sequence modeling techniques. Results motivate us to explore and optimize the proposed approach further to enhance its performance in generative audio tasks.

### v2

@@ Abstract
Recent advancements in deep learning have led to widespread use of techniques for audio content generation, notably employing Denoising Diffusion Probabilistic Models (DDPM) across various tasks. Among these, Foley Sound Synthesis is of particular interest for its role in applications for the creation of multimedia content. Given the temporal-dependent nature of sound, it is crucial to design generative models that can effectively handle the sequential modeling of audio samples. Selective State-Space Models (SSMs) have recently been proposed as a valid alternative to previously proposed techniques, demonstrating competitive performance with lower computational complexity. In this paper, we introduce MambaFoley, a diffusion-based model that, to the best of our knowledge, is the first to leverage the recently proposed SSM known as Mamba for the Foley sound generation task. To evaluate the effectiveness of the proposed method, we compare it with a state-of-the-art Foley sound generative model using both objective and subjective analyses.
@@
Foley Sound Synthesis, State-Space Models, Generative Sound Synthesis, Denoising Diffusion Probabilistic Models
@@ I Introduction
In recent years, audio generation using deep-learning generative techniques has made relevant progress. This was possible primarily by the adoption of generative models denoted as Denoising Diffusion Probabilistic Models (DDPMs) [1], which have been applied to a wide variety of audio applications such as Text-to-Audio generation [2, 3, 4], sound field reconstruction [5], among others. Generated audio samples from generative models are so realistic that using them to replace real datasets for training deep learning models has been explored across various domains [6, 7].
@@
Generating audio waveforms with mid and long-term temporal coherence has always been a challenge for learning-based models. In fact, although using raw waveforms in audio generative modeling can significantly improve the quality of generated audio content, they also pose challenges due to the high-dimensional input signal, which includes tens of thousands of samples. Several deep learning models have been proposed to address the challenge of modeling raw waveform sequences, each with its own drawbacks [17, 18, 19, 20]. RNNs based on LSTMs [17] are not parallelizable, which limits their computational efficiency. CNNs [18] are constrained by their receptive field, which limits their ability to model the global coherence of the waveform. Transformers and, in general, attention-based methods [19, 20], while effective for modeling sequential data, are impractical for very long sequences due to their quadratic complexity with respect to window length. Recently, structured State-Space Models (SSMs) [21, 22] have been proposed as a new class of architectures tailored to sequence modeling, combining characteristics of both CNNs and RNNs. Architectures such as SaShiMi [23] and especially Mamba [24, 25] outperformed previous state-of-the-art techniques over waveform modeling while scaling linearly in complexity with respect to sequence length. As expected, Mamba models were readily applied to a diverse variety of audio signal processing tasks such as audio representation learning [26], deepfake detection [27], and speech separation [28]. Motivated by previous studies, we explored how to integrate selective SSMs into the Foley sound generation task.
In this paper, we introduce MambaFoley, a method for Foley sound generation that leverages the SSM Mamba to model raw audio. To the best of our knowledge, this is the first application of Mamba in Foley sound generation. The proposed model employs the diffusion architecture DAG [29] and integrates Mamba in the bottleneck to improve its sequence modeling capabilities. Additionally, we enable temporal control of the generated Foley sounds through the BlockFiLM conditioning method introduced in [14]. Fig. 1 shows a schematic overview of the MambaFoley model. To validate the effectiveness of our proposed model, we conduct a series of experiments comparing it with the model proposed in [14] and an Attention-based baseline. We evaluate performance using both objective metrics and subjective evaluations. The rest of the paper is organized as follows: Sec. II provides the reader with the necessary background on SSMs. In Sec. III, we present design choices of the MambaFoley model. Sec. IV describes how the experiments were organized, while Sec. V presents the results, highlighting the effectiveness of the proposed technique in comparison to state-of-the-art methods. In Sec. VI we draw some conclusions. Code and listening examples are publicly available online [^1] https://furiocolombo.github.io/mamba-foley.
@@ III-A Problem formulation
The goal of MambaFoley is to generate raw audio waveforms that represent specific sound classes (e.g. dog, gunshot, etc.) while also following a specified temporal pattern (e.g. two dog barks separated by one second of silence).
More formally, let $\mathbf{x}\in\mathbb{R}^{N}$ be a discrete audio waveform whose content corresponds to the category indicated by the class $c\in\{0,\ldots,C-1\}$ of length $N$ samples We want to control the temporal evolution of $\mathbf{x}$ using $\boldsymbol{\tau}\in\mathbb{R}^{I}$, which can be defined using different representations.
MambaFoley then learns the function $\mathcal{M}(\cdot)$ such that $\mathbf{x}=\mathcal{M}(c,\boldsymbol{\tau}).$ Following the approach proposed in [14], we condition the temporal evolution of the generated audio using the Root-Mean-Square (RMS) of a waveform signal that contains an audio excerpt of the target category with the desired temporal evolution. Specifically, in our scenario we can define:
@@ III-B Network Architecture
MambaFoley is a diffusion-based model [1], where the target audio waveform $\mathbf{x}$ is progressively corrupted with Gaussian noise throughout the forward process. In the backward process, the original waveform is iteratively denoised using a U-Net architecture, whose parameters are updated based on the $\ell_{2}$ norm between the predicted and ground truth noise $\boldsymbol{\epsilon}$. The overall architecture of MambaFoley is illustrated in Fig. 2.
We select the DAG architecture [29] as the U-Net backbone of our model, due to its proven effectiveness in generating high-resolution raw audio. We introduce some changes to DAG, specifically regarding the conditioning and bottleneck of the model.
Similar to DAG, we refer to the building blocks of MambaFoley as GBlocks, to which we apply significant modifications. GBlocks consist of sequences of 1D-convolutional layers in the downsampling encoder part, or transposed 1D-convolutions in the upsampling decoder. Fig. 3 (a) shows the architecture of GBlocks.
@@
In the bottleneck of the model, between the encoder and decoder parts of the U-Net, we introduce a bidirectional Mamba layer [28, 31] to exploit the capabilities of SSMs in processing temporal audio sequences. This layer consists of two parallel Mamba blocks, as defined in Sec. II, one processing the input sequence and the other one processing its reversed version. Fig. 3 (b) shows the bidirectional Mamba bottleneck. The outputs of the two parallel Mamba blocks are each normalized using Root Mean Square Layer Normalization (RMSNorm) [32], chosen for its computational efficiency. The normalized outputs are then summed with the input (and reversed) signal residuals. Finally, the outputs of the two SSMs are concatenated and processed by a single linear layer. We use state dimension D=$16$, expansion factor $4$ and local convolution width $4$ for the Mamba layers.
@@ IV-C Training details
All models are trained for $500$ epochs using the Adam [36] optimizer with a learning rate of 1e-4. To make the training more robust, we randomly train the network unconditionally with a $0.1$ probability. We adopt a variance-preserving cosine scheduler, as proposed in  [37]. For conditional sampling, we employ the DDPM-like discretization of SDE [38] with classifier-free guidance [39] to enhance sample diversity and quality. For the temporal event features, we use a window length of $W=512$ samples and a hop size of $H=128$ samples, which correspond to approximately $24~{}\mathrm{ms}$ and $6~{}\mathrm{ms}$, respectively, given the sampling rate of $22.05$ kHz. For a fair comparison with T-Foley [14], we also exclude the SneezeCough class from the listening test.
@@
Figure 4: Example spectrograms (c-e) of samples generated using different models. The category conditioning corresponds to the class gunshot, while the temporal conditioning is provided via the RMS shown in (a) and computed over the ground truth shown in (b).
@@ V Results
In this section, we present results demonstrating the effectiveness of the proposed MambaFoley model compared to T-Foley and AttentionFoley baselines, both in terms of objective and subjective evaluation metrics. Before presenting the analyses computed on the entire dataset, we examine a representative spectrogram generated from the waveforms obtained through each of the considered models, shown in Fig. 4. The STFT for all spectrograms is computed with a Hann window of length $512$, hop size of $128$ samples and a FFT of $512$ points. Fig. 4(b) contains an example of the target sound, corresponding to a gun shot, from which we extract the RMS, shown in Fig.4 (a). The RMS is then used to condition the temporal evolution of the network. As observed from the spectrograms of all the considered models in Fig. 4(c) (MambaFoley), (d) (T-Foley), and (e) (AttentionFoley), each model appears to accurately reproduce the desired temporal conditioning. However, MambaFoley seems to reconstruct more content at higher frequencies.
@@ V-A Objective Evaluation Results
Real data - - - 0 15.20 3.31
[…]
Table I reports the results for E-L1 and FAD for the three different feature extractors considered in this study. It also includes the number of parameters for each network, required floating-point operations per second (FLOPs) and the generation time for a single audio sample. To facilitate the analysis, we highlight the best result in each category in bold. The top row presents the lower bound on the achievable error, represented by the FAD and E-L1 scores computed on real data. Although the lower bound of the E-L1 distance is clearly 0, the FAD calculated between the embeddings of the test and train datasets for real data is greater than zero. This indicates the inherent differences between real audio samples. This finding helps to estimate the lower bound for FAD in our task, where our goal is to generate as realistic audio samples as possible.
Table I shows that MambaFoley achieves faster inference and outperforms both T-Foley and AttentionFoley in terms of FAD. The models demonstrate comparable E-L1 performance, likely due to the influence of the conditioning blocks over the bottleneck, which help maintain consistent temporal alignment.
@@ VI Conclusion
In this paper, we introduced MambaFoley, a Denoising Diffusion Probabilistic model for Foley sound generation exploiting the power of selective State-Space models. Specifically, we have applied Mamba in the bottleneck of a diffusion model, a SSM that has proved to obtain notable performance in audio sequence modeling, while mantaining computational complexity. Through a series of experiments that include both objective and subjective results, we demonstrate the effectiveness of the proposed approach by comparing it with baseline methods using various sequence modeling techniques. Results motivate us to explore and optimize the proposed approach further to enhance its performance in generative audio tasks.